OpenStudy (anonymous):

Find the angle between the lines 3x+2y-4=0 and 4x-y+6=0.

OpenStudy (anonymous):

do you know how to find the angle of inclination of lines?

OpenStudy (anonymous):

\[\theta = \arctan \frac{\left(-\frac32\right)- 4}{1 + \left(-\frac32\right)\cdot 4}\]